You have a 4 ohm resistor and a 6 ohm resistor connected in series with a 12.0 volt battery. What is the voltage drop across the 6 ohm resistor (in volts)?
Answer #1

here given

4 \Omega resistor and a 6 \Omega resistor connected in series

with a 12.0 volt battery = V

- - 465 120

4 \Omega and 6 \Omega are connected in series

Req = R1 + R2

= 4 + 6

Req = 10 \Omega

using Ohm's Law V = i Req

i = V / Req

= 12 / 10

i = 1.2 A

the voltage drop across the 6 \Omega resistor is say V ' = ?

V ' = i R6 ohm

= 1.2 x 6

V ' = 7.2 volts

the voltage drop across the 6 \Omega resistor is say V ' = 7.2 volts